OpenStudy (anonymous):

Solve: 5x − 10 + 12 = 22

mathslover (mathslover):

\[\LARGE{5x-10+12=22}\] Ok! can you tell me ... what is 12 - 10 ?

OpenStudy (anonymous):

5x=20 x=4

OpenStudy (anonymous):

thanks all

OpenStudy (anonymous):

Solve: −2(x − 1) = −16

OpenStudy (anonymous):

2(x-1)=16 (x-1)=16/2=8 x=8+1 x=9